# Investigating MAXI J1752--457 with NuSTAR in the aftermath of a superburst

### v1

@@ 4 Discussion and Conclusions
This work was partially supported under NASA Grant No. 80NSSC25K7455 and made use of data from the NuSTAR mission, a project led by the California Institute of Technology, managed by the Jet Propulsion Laboratory, and funded by the National Aeronautics and Space Administration. We thank the NuSTAR Operations, Software, and Calibration teams for support with the execution and analysis of these observations. This research has made use of the NuSTAR Data Analysis Software (NuSTARDAS), jointly developed by the ASI Science Data Center (ASDC, Italy) and the California Institute of Technology (USA). This research has also made use of the MAXI data provided by RIKEN, JAXA and the MAXI team.

### v2

@@ V Conclusions
This work was partially supported under NASA Grant No. 80NSSC25K7455 and made use of data from the NuSTAR mission, a project led by the California Institute of Technology, managed by the Jet Propulsion Laboratory, and funded by the National Aeronautics and Space Administration. This work was also supported by Grants-in-Aid for Scientific Research 21K03620 (H.N.) from the Ministry of Education, Culture, Sports, Science and Technology (MEXT) of Japan. We thank the NuSTAR Operations, Software, and Calibration teams for support with the execution and analysis of these observations. This research has made use of the NuSTAR Data Analysis Software (NuSTARDAS), jointly developed by the ASI Science Data Center (ASDC, Italy) and the California Institute of Technology (USA). This research has also made use of the MAXI data provided by RIKEN, JAXA and the MAXI team. We would like to acknowledge the Einstein Probe team, with particular thanks to Weimin Yuan and Dongyue Li, for their assistance in determining the pre-superburst flux of the source. Finally, we would like to thank the anonymous reviewer for their feedback which improved the quality of this paper.